i wants to buy flowers for my friend with $27. daisies are $1 each and roses are $3 each. but i buys 3 more daisies than roses. how much did the roses cost?

Answer by mananth(12270)

let number of roses be x each $3
daisies bought will be x+3 at $1
..
total cost = $27
...
3x+1(x+3)=27
3x+x+3=27
4x+3=27
-3
4x=27-3
4x=24
/4
x= 6 roses.
6 roses cost $18
...
m.ananth@hotmail.ca